Marc Horowitz

24.151

February 14, 1992

A) The sentences in (13) and (14) do not have the same cognitive significance. Although they may have the same truth value and truth conditions, given facts from astronomical observations, the information content is different. Sentence (14) is a tautology: ‘If P then P’ is always true. Sentence (13) is much more complicated. To evaluate the truth of this sentence, we must find the relationship between the morning and evening stars, if there is one. It is not intuitively obvious that the referents of ‘morning star’ and ‘evening star’ are the same. The first sentence can be proven or disproven merely by using logic; it exists regardless of the nature of our physical universe. To verify the second requires some real knowledge.

B) According to the ‘Fido’-Fido theory, the proposition ‘Aphrodite = Eros’ is true. Since the referents do not exist, they can be viewed as being the same. This results in a problem for ‘Fido’-Fido, since we can substitute as co-referentials expressions which we observe as different. For example, if I have a story about Eros, I should be able to substitute ‘Aphrodite’ for ‘Eros’ and still have a cognitively equivalent story. But, this is clearly not the case to anyone reading the two stories.